5. There are two barrels of gasoline in the warehouse. There are 21 kilograms in barrel A and 5 kilograms in barrel B. After adding the same quality gasoline, the gasoline in barrel A is 3 times that in barrel B. Now there are two barrels A and B. How many kilograms of gasoline are in the barrel?


1
Expert's answer
2021-08-10T06:42:42-0400

Suppose that we add x kilograms in barrel A and 5 kilograms in barrel B. Then


"21+x=3(5+x)"

"21+x=15+3x"

"2x=6"

"x=3"

"21+3=24"

"5+3=8"

There are 24 kilograms in barrel A and 8 kilograms in barrel B.